# Spin-1/2 kagome Heisenberg antiferromagnet: Machine learning discovery of the spinon pair density wave ground state

### v1

@@ Abstract
Spin-1/2 kagome antiferromagnet (AFM) is one of the most studied models in frustrated magnetism since it is a promising candidate to host exotic spin liquid states. However, despite numerous studies using both analytical and numerical approaches, the nature of the ground state and low-energy excitations in this system remain elusive. This is related to the difficulty in determining the spin gap in various calculations. We present the results of our investigation of the Kagome AFM using the recently developed group equivariant convolutional neural networks – an advanced machine learning technique for studying strongly frustrated models. The approach, combined with variational Monte Carlo, introduces significant improvement of the achievable results accuracy in comparison with approaches based on other neural network architectures that lack generalization quality for frustrated spin systems. Contrary to the results obtained previously with various methods, that predicted $Z_{2}$ or U(1) Dirac spin liquid states, our results strongly indicate that the ground state of the kagome lattice antiferromagnet is a spinon pair density wave that does not break time-reversal symmetry or any of the lattice symmetries. The found state appears due to the spinon Cooper pairing instability close to two Dirac points in the spinon energy spectrum and resembles the pair density wave state studied previously in the context of underdoped cuprate superconductors in connection with the pseudogap phase. The state has significantly lower energy than the lowest energy states found by the SU(2) symmetric density matrix renormalization group calculations and other methods.
@@ I Introduction
Among many numerical approaches applied to study kagome antiferromagnet properties are exact diagonalization (ED),Lauchli ; Wiatek ; Changlani2 density matrix renormalization group (DMRG),Jiang ; Yan ; Depenbrock ; Jiang2 ; Nishimoto ; Bauer ; Gong ; He tensor network methodsEvenbly ; Xie ; Mei ; Niu ; Liao ; Haghshenas ; Jiang3 and variational Monte Carlo (VMC).Ran ; Iqbal ; Clark2 ; Iqbal2 ; Tay ; Hu ; Hu2 ; Iqbal3 ; Westerhout ; Fu ; Kochkov ; Yang Different studies suggested so far a variety of candidate ground states. In addition to initially proposed valence-bond crystal state with a 36-site unit cell,Evenbly ; Marston ; Zeng ; Nikolic ; Singh1 ; Singh2 ; Poilblanc ; Schwandt ; Poilblanc2 ; Hwang ; Poilblanc3 majority of recent DMRG and VMC studies find spin-liquid ground states. Whilst SU(2) symmetric DMRG calculations obtain a fully gapped (nonchiral) $\mathbb{Z}_{2}$ topological ground state,Depenbrock DMRG calculations supplemented with additional flux insertion indicate a gapless algebraic $U(1)$ Dirac spin liquid ground stateHe found also within the projected fermionic wave-function approaches Ran ; Iqbal ; Iqbal2 ; Iqbal3 ; Hermele ; Ma ; Iqbal4 ; Iqbal5 and the tensor-network states formalism.Liao ; Lee
Difficulty in finding quantum spin liquids in realistic microscopic models is largely due to lack of adequate unbiased computational methods. Unbiased quantum Monte Carlo simulations face still unsolved sign problemPan ; Troyer while tensor network and many variational approaches are often biased towards particular kinds of states. Also, although DMRG method White ; White2 ; Schollwock ; Schollwock2 is known to be very accurate method for calculating properties of one-dimensional systems, difficulty of obtaining reliable results with DMRG greatly increases for $D$-dimensional systems with $D>1$. In addition, matrix product states within DMRG algorithm are biased towards low-entangled wave-functions.Chan Gapped low-entangled states are therefore more favored within DMRG algorithm than highly entangled gapless phases.
@@
In our calculations we employ advanced neural-networks based machine learning (ML) techniques to study properties of the kagome lattice antiferromagnet since these novel unbiased ML techniques have so far shown great promise and potential for studying various many-body problems. An important advantage of the ML approach is suitability for implementation on graphics processing units (GPUs) that have enhanced computational capability. Namely, GPUs can perform multiple simultaneous computations and provide significant computational process speedup. Additional benefit is also powerful ML optimization libraries like NetKetCarleo2 ; Vicentini2 and TensorFlowAbadi libraries.
Neural network quantum states (NQS) have recently been recognized as powerful variational wave-function ansätze that allow accurate representation of various quantum many-body states. Carleo and Troyer proposed an NQS ansatz based on the restricted Boltzmann machine neural networkCarleo that was initially introduced for quantum spin systems without geometrical frustration. The applicability of various NQS ansätze has later been extended to many complex challenging many-body problems for which the exact solutions are still not known. Among many examples are frustrated spin systems,Cai ; Liang ; Choo3 ; Ferrari ; Sharir ; Hibat-Allah ; Westerhout ; Szabo ; Nomura3 ; Astrakhantsev ; Nomura2 ; Bukov ; Fu ; Kochkov ; Yang ; Roth ; Roth2 fermionic modelsNomura ; Cai ; Luo ; Pfau ; Choo ; Stokes ; Cassella and topological phases of matter.Changlani ; Clark ; Deng ; Glasser ; Kaubruegger ; Lu ; Huang ; Vieijra ; Vieijra2 In addition to the initial NQS based calculations that considered mainly ground-state properties, neural network based numerical methods have also been further developed to study excited energy states,Nomura3 ; Nomura2 ; Nomura ; Roth ; Roth2 ; Vieijra ; Vieijra2 ; Duric ; Choo2 time dynamics of many-body quantum systems,Schmitt ; Gutierrez ; Hofmann open quantum systemsNagy ; Hartmann ; Vicentini ; Yoshioka ; Luo4 ; Reh2 and thermal quantum states at finite temperatures.Irikura ; Nomura4
@@
We study properties of the spin-1/2 kagome lattice antiferromagnet by representing the ground and low-lying excited energy states by group equivariant convolutional neural network (GCNN) ansätze equivariant over the full space group composed of transformations that leave the kagome lattice unchanged (lattice translations and dihedral D(6) point group transformations). In addition of being very expressive ansätze GCNN also possess greater ability to learn from a limited number of samples, that is greater generalization quality, in comparison with other neural network architectures. This improved generalization quality allows more efficient neural network training yielding to significant improvement of the achievable results accuracy within the GCNN approach combined with VMC. Unlike density matrix renormalization group and other variational methods including tensor network approaches, the GCNN and VMC approach does not exhibit significant bias towards any particular kind of states and therefore provides an important step in clarifying the kagome antiferromagnet properties and a reference to check validity of the results obtained with other methods.
Our numerical calculations were performed on NVIDIA A100 80GB GPUs and using NetKet,Carleo2 ; Vicentini2 JAX,Frostig FLAX,Heek and OPTAXHessel libraries. Contrary to the results obtained previously with other methods, we find that the kagome lattice antiferromagnet ground state is a spinon pair density wave (PDW) state that does not break time-reversal symmetry or any of the lattice symmetries. The state appears due to the spinon Cooper pairing instability close to two Dirac nodes of the underlying $U(1)$ Dirac spin liquid found with recent flux insertion supplemented DMRG approachHe and also with projected fermionic wave-function approachesRan ; Iqbal ; Iqbal2 ; Iqbal3 ; Hermele ; Ma ; Iqbal4 ; Iqbal5 and within the tensor-network states formalism.Liao ; Lee The found spinon PDW resembles the pair density wave state studied previously in the context of underdoped cuprate superconductors in connection with pseudogap phaseAgterberg ; Lee3 ; Lee4 ; Xu ; Chakraborty and has significantly lower energy than the ground state energy values obtained by other methods. In particular, for the studied system sizes we find that the PDW energy has $\approx 1.78\%$ lower value than the ground state energy found by SU(2) symmetric density matrix renormalization group calculations.Depenbrock
The paper is organized as follows. In Sec. II we describe the GCNN wave-function ansätze. In Sec. III we explain wave-function optimization procedure. Our results for the ground state properties and low-energy excited states are presented in Sec. IV. We summarize our results, draw conclusions and discuss future research directions in the last section, Sec. V.
II Wave-function representation with group convolutional neural networks

### v2

@@ Abstract
Spin-1/2 kagome antiferromagnet (AFM) is one of the most studied models in frustrated magnetism since it is a promising candidate to host exotic spin liquid states. However, despite numerous studies using both analytical and numerical approaches, the nature of the ground state and low-energy excitations in this system remain elusive. This is related to the difficulty in determining the spin gap in various calculations. We present the results of our investigation of the Kagome AFM using the recently developed group equivariant convolutional neural networks – a novel machine learning technique for studying strongly frustrated models. The approach, combined with variational Monte Carlo, introduces significant improvement of the achievable results accuracy for frustrated spin systems in comparison with approaches based on other neural network architectures. Contrary to the results obtained previously with various methods, that predicted $Z_{2}$ or U(1) Dirac spin liquid states, our results strongly indicate that the ground state of the kagome lattice antiferromagnet is a spinon pair density wave that does not break time-reversal symmetry or any of the lattice symmetries. The found state appears due to the spinon Cooper pairing instability close to two Dirac points in the spinon energy spectrum and resembles the pair density wave state studied previously in the context of underdoped cuprate superconductors in connection with the pseudogap phase. The state has significantly lower energy than the lowest energy states found by the SU(2) symmetric density matrix renormalization group calculations and other methods.
@@ I Introduction
Among many numerical approaches applied to study kagome antiferromagnet properties are exact diagonalization (ED),Lauchli ; Wiatek ; Changlani2 density matrix renormalization group (DMRG),Jiang ; Yan ; Depenbrock ; Jiang2 ; Nishimoto ; Bauer ; Gong ; He tensor network methodsEvenbly ; Xie ; Mei ; Niu ; Liao ; Haghshenas ; Jiang3 and variational Monte Carlo (VMC).Ran ; Iqbal ; Clark2 ; Iqbal2 ; Tay ; Hu ; Hu2 ; Iqbal3 ; Westerhout ; Fu ; Kochkov ; Yang Different studies suggested so far a variety of candidate ground states. Aside from an initially proposed valence-bond crystal state with a 36-site unit cell,Evenbly ; Marston ; Zeng ; Nikolic ; Singh1 ; Singh2 ; Poilblanc ; Schwandt ; Poilblanc2 ; Hwang ; Poilblanc3 majority of recent DMRG and VMC studies find spin-liquid ground states. Whilst SU(2) symmetric DMRG calculations obtain a fully gapped (nonchiral) $\mathbb{Z}_{2}$ topological ground state,Depenbrock DMRG calculations supplemented with additional flux insertion indicate a gapless algebraic $U(1)$ Dirac spin liquid ground state.He The latter is also found in the projected fermionic wave-function approaches Ran ; Iqbal ; Iqbal2 ; Iqbal3 ; Hermele ; Ma ; Iqbal4 ; Iqbal5 and the tensor-network states formalism.Liao ; Lee
Difficulty in finding quantum spin liquids in realistic microscopic models is largely due to lack of adequate computational methods. Whilst quantum Monte Carlo simulations face still unsolved sign problem,Pan ; Troyer tensor network and many variational approaches often favor particular kinds of states. Also, although DMRG method White ; White2 ; Schollwock ; Schollwock2 is known to be very accurate method for calculating properties of one-dimensional systems, difficulty of obtaining reliable results with DMRG greatly increases for $D$-dimensional systems with $D>1$. In addition, matrix product states within DMRG algorithm are more suitable to study gapped low-entangled states than highly entangled gapless phases.Chan
@@
In our calculations we employ novel neural-networks based machine learning (ML) techniques to study properties of the kagome lattice antiferromagnet since these techniques have so far shown great promise and potential for studying various many-body problems. An important advantage of the ML approach is suitability for implementation on graphics processing units (GPUs) that have enhanced computational capability. Namely, GPUs can perform multiple simultaneous computations and provide significant computational process speedup. Additional benefit is also powerful ML optimization libraries like NetKetCarleo2 ; Vicentini2 and TensorFlowAbadi libraries.
Neural network quantum states (NQS) have recently been recognized as powerful variational wave-function ansätze that allow accurate representation of various quantum many-body states. Carleo and Troyer proposed an NQS ansatz based on the restricted Boltzmann machine neural networkCarleo and demonstrated it for quantum spin systems without geometrical frustration. Subsequently, NQS ansätze have later been extended to many complex many-body problems for which the exact solutions are still not known, such as frustrated spin systems,Cai ; Liang ; Choo3 ; Ferrari ; Sharir ; Hibat-Allah ; Westerhout ; Szabo ; Nomura3 ; Astrakhantsev ; Nomura2 ; Bukov ; Fu ; Kochkov ; Yang ; Roth ; Roth2 ; Beck ; Viteritti fermionic modelsNomura ; Cai ; Luo ; Pfau ; Choo ; Stokes ; Cassella and topological phases of matter.Changlani ; Clark ; Deng ; Glasser ; Kaubruegger ; Lu ; Huang ; Vieijra ; Vieijra2 In addition to their initial application to determine ground state properties, NQS based methods have been further developed to study excited energy states,Nomura3 ; Nomura2 ; Nomura ; Roth ; Roth2 ; Vieijra ; Vieijra2 ; Duric ; Choo2 time dynamics of many-body quantum systems,Schmitt ; Gutierrez ; Hofmann open quantum systemsNagy ; Hartmann ; Vicentini ; Yoshioka ; Luo4 ; Reh2 and thermal quantum states at finite temperatures.Irikura ; Nomura4
@@
We study the spin-1/2 kagome lattice antiferromagnet by representing the ground and low-lying excited energy states by group equivariant convolutional neural network (GCNN) ansätze equivariant over the full space group composed of transformations that leave the kagome lattice unchanged (lattice translations and dihedral D(6) point group transformations). GCNNs are very expressive ansätze that can efficiently learn from a limited number of samples. This yields to significant improvement of the achievable results accuracy within the GCNN approach combined with VMC. The GCNN and VMC approach is therefore a powerful new numerical approach for clarifying the kagome antiferromagnet properties that can provide a reference to check validity of the results obtained with other methods.
Our numerical calculations were performed using computational speedup on GPUs and NetKet,Carleo2 ; Vicentini2 JAX,Frostig FLAX,Heek and OPTAXHessel libraries. Contrary to the results obtained previously with other methods, we find that the kagome lattice antiferromagnet ground state is a spinon pair density wave (PDW) state that does not break time-reversal symmetry or any of the lattice symmetries. The state appears due to the spinon Cooper pairing instability close to two Dirac nodes of the underlying $U(1)$ Dirac spin liquid found with recent flux insertion supplemented DMRG approachHe and also with projected fermionic wave-function approachesRan ; Iqbal ; Iqbal2 ; Iqbal3 ; Hermele ; Ma ; Iqbal4 ; Iqbal5 and within the tensor-network states formalism.Liao ; Lee The found spinon PDW resembles the pair density wave state studied previously in the context of underdoped cuprate superconductors in connection with pseudogap phaseAgterberg ; Lee3 ; Lee4 ; Xu ; Chakraborty and has significantly lower energy than the ground state energy values obtained by other methods. In particular, for the studied system sizes we find that the PDW energy has $\approx 1.78\%$ lower value than the ground state energy found by SU(2) symmetric density matrix renormalization group calculations.Depenbrock
The paper is organized as follows. In Sec. II we describe the GCNN wave-function ansätze and explain wave-function optimization procedure. Our results for the ground state properties and low-energy excited states are presented in Sec. III. Description of the found spinon pair density wave ground state and possible induced secondary orders within Abrikosov fermion parton construction is presented in Sec. IV. We summarize our results, draw conclusions and discuss future research directions in the last section, Sec. V.
II GCNN wave-function representation and training scheme for the wave-function optimization